OpenStudy (anonymous):

Solve: 4x + 16 + 40 = 160

OpenStudy (anonymous):

4x = 160 - 40 - 16 4x = 120-16 4x = 104 x = 26

OpenStudy (anonymous):

4x=104 x=26 answer

OpenStudy (anonymous):

4x+16+40=160 becomes 4x+56=160 -56 -56 ------------- 4x=104 -- --- 4 4 -------- x= ?